OpenStudy (anonymous):

Consider the sequence 2,6,18,54,... Let n = the term number in the sequence. Let A(n) = the value of the nth term of the sequence. What is the comon ratio of the sequence.? Complete each statement. a. A(1) = 2 = 2 x 3^? b. A(2) = 6 = 2 x 3 = 2 x 3^? c. A(3) = 18 = 2 x 3 x 3 = 2 x 3^? d. A(4) = 54 = 2 x 3 x 3 x 3 = 2 x 3^? Thanks! (: Medals will be given lol.

OpenStudy (anonymous):

The x's are mutiplication and you dont know the exponets.

OpenStudy (aravindg):

if a,b,c form a GP Then common ratio is given by r=b/a=c/b

OpenStudy (aravindg):

now can u tell the common ratio here?

OpenStudy (anonymous):

I know the common ratio is 3, I don't know how to do the bottom./:

OpenStudy (aravindg):

nth term of a GP is given by ar^(n-1) where a is first term and r is common ratio ..i hope u can do it now :)

OpenStudy (anonymous):

Oh okiee. Thanks.

OpenStudy (aravindg):

yw :)